hi can you please help me with this question...it says:
substitute a=-12 and b=-2 to find the value of this expression
2(a-5b)
thank you

Found 2 solutions by queenbuc, dolphingurl:
Answer by queenbuc(2)

To solve substitute the letters for the numbers.
Everywhere you see a, put in -12
Everywhere you see b, put in a -2
2(a-5b)= ?
2(-12- 5(-2))=
2(-12--10)=
2(-12+10)=
2(-2)=-4

Answer by dolphingurl(4)

To solve substitute the letters for the numbers.
Everywhere you see a, put in -12
Everywhere you see b, put in a -2
2(a-5b)= ?
2(-12- 5(-2))=
2(-12--10)=
2(-12+10)=
2(-2)=-4